You are working for a window manufacturing company. The company has just landed a new contract for homes being built in Alaska. There is some concern about the windows surviving in an area where fierce hailstorms occur with some regularity. Because of strong horizontal winds, the hailstones often fall at an angle of θ = 45.0° to the ground. Measurements have been made during such winds with the following typical results: hailstones strike vertical surfaces at an average rate of 2,300 per minute per square meter of vertical surface. The hailstones have an average mass of m = 5.00 and an average speed of v = 7.70 m/s. For your report to your supervisor, determine the following on a window of area A = 0.650 m2 due to the hailstones. (Assume the collisions are elastic.)

(a) the average force (in N)

(b) the average pressure (in Pa)
Answer #1

Change in momentum of 1 stone after collison along x=2mvcos45

=2*5*10^-3*7.7(cos45)=0.054

Assuming m=5g

Force=change in momentum/time

Force=2300*0.054*.65/60=1.36N

Pressure=Force/Area

=1.36/0.65

=2.087Pa